# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

### v1

@@ III Results
The excitation-evolution-emission events occur in three stages. Excitation by entangled photon pairs creates a tailored two-exciton wavepacket. The two-exciton wavepacket or density operator populations undergo transport, leading to a redistribution of populations over a finite waiting time. This redistributed population serves as the initial configuration for photon emission events. The cascaded two-photon emission process proceeds via two one-photon emissions interleaved by exciton transport in the one-exciton manifold. The signal can be expressed in a compact superoperator form
@@
where the superoperator for filtered detection
@@
and the bare two-photon signal
@@
is defined using the following superoperators corresponding to emission events
@@
The averaging has been carried out over the two-exciton density operator following transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$. Under a set of physically motivated approximations, the averaging can be carried out over each of these three terms independently. We also define the entangled photon driven excitation superoperator for future convenience, given by
$O(t)=T_{R}^{{\dagger}}(\tau_{2},\tau_{3})T_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
@@
In the above, we assume that the excitation and emission events are not temporally overlapping, which results in the signal in Eq. 9 being factorized in terms of three principal contributions: first, preparation, involving exciton-photon (driving fields) interactions that lead to the preparation of the two-exciton density operator; second, evolution, describing two-exciton transport over a parametric time interval; and third, detection, describing exciton-photon (detection fields) interactions that yield the time-frequency filtered two-photon coincidence. In subsequent sections, we systematically describe all three contributions.
@@ III.1 Entangled photon-pair excitation
The quantity of interest at the preparation stage, the two-exciton population distribution, can be defined using a projection operator acting on the diagonal matrix elements of the phonon-averaged exciton density operator, expressed as: $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]_{\rho_{0}}$. Here, a factorizable initial condition between the driving, the detection field modes, and the exciton density operator is assumed. We perturbatively expand the combined photon-exciton-phonon density operator to the fourth order in $H_{\text{int}}(t)$, and partially trace over the phonon modes to obtain the desired expression for the observable (derivation detailed in Appendix B) as follows:
@@
$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})$
@@
$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (14)
It is a sum of five distinct terms that represent interfering exciton pathways (see Liouville space Feynman diagrams presented in Fig. 2 for more insights). It includes contributions from both one- and two-exciton coherence and population terms of the exciton density operator. The first term represents the two-exciton coherence-mediated pathway, which is susceptible to inter-manifold dephasing. The second and fourth terms describe pathways susceptible to one-exciton population transport. The third and fifth terms describe pathways susceptible to intra-manifold dephasing. The convolutional form is particularly advantageous for introducing the field correlation functions, given in Eq. 6-Eq. II. The field correlation functions can be manipulated to selectively suppress and amplify specific exciton resonances. Next, we illustrate these capabilities via numerical simulations for both degenerate and nondegenerate entangled photon pairs.
@@
In this section, we monitor the final population distribution following the excitation via spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. We had carried out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig. 3. The simulation presented in the left column uses an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, and acts as a reference case. Variations in the temporal width of the SPDC pump ($\tau_{0}=50\,\text{fs}$) and the entanglement time parameter ($T_{\text{ent}}=60\,\text{fs}$) are presented in the middle and the right column, respectively. We note that target states near the middle of the energy ladder exhibited broader distributions around the target in all three cases. In other words, we find a lower degree of selectivity for target states in the middle rung of the two-exciton manifold and higher selectivity for those belonging to the lower rung. Altering the width of SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways. Altering the entanglement time in the right column preserved the selectivity for the majority of states. Across all plots, we notice a trend: some two-exciton states remain preferentially populated across all parameter regimes. This feature have likely emerged from an intricate interplay of exciton transport parameters leading to the absence of specific relaxation channels.
@@
Now, we focus on two specific two-exciton states, $f_{07}$ and $f_{83}$, among the full suite of target states previously monitored, and examine the possibility of their excitation via a chosen pair one-exciton states, selected by setting $\omega_{1}=E_{07}$ and $\omega_{2}=E_{09}$. The mediating states are delocalized over sites and prone to ultrafast transport. The results are displayed in the left column of Fig. 4, and Fig. 5, respectively. The entanglement time and width of SPDC pump are set at $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$, for the upper panel (similar to the reference case in Fig. 3). The width is varied to $\tau_{0}=50\,\mathrm{fs}$ in the middle panel, and the entanglement time is varied to $T_{\mathrm{ent}}=40\,\mathrm{fs}$ in the bottom panel. The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy. The increase in the temporal width results in a significant loss of selectivity as the number of mediating pathways with comparable spectral weights grows. The variation of the entanglement time didn’t result in any major changes, except for the tailoring of the relative populations. The latter, however, will have consequences as the population redistributes over time, as will be shown in the next section.
@@ III.2 Transport in the two-exciton manifold
The created population distributions undergo time evolution governed by the parameters of the two-exciton transport matrix. Here, the following questions are worth investigating: What is the nature of the population evolution for a given state distribution? How do two initially created two-exciton population distributions undergo redistribution? And how does the nature of redistribution kinetics vary with driving field parameters?
The kinetic equation that governs the dissipative time evolution of exciton populations, i.e., exciton transport is given by
@@
where the transport rate matrices, $K_{ff,f^{\prime}f^{\prime}}$ follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$. The initial two-exciton population distribution is obtained via the Eq. III.1 which is propagated using the expression: $\rho_{ff}(t_{2})=G_{ff,f^{\prime}f^{\prime}}^{(N)}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$, where the $G_{f^{\prime}f^{\prime},ff}^{(N)}(\tau)$ is given by first term in Eq. II. The following four sets of simulations will help us address the questions posed above.
@@
We track the differential population evolution following the excitation targeted at $f_{07}$ (right column in Fig. 4) and $f_{83}$ (right column in Fig. 5), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively. Each of the three panels, upper, middle and bottom, represents simulation using different photonic parameters. The parameters were listed in the preceding section III.1.2 where we discussed the reference simulation. We observe that the evolved distributions at short times retained the memory of prepared distribution in Fig. 4, but gets washed out in Fig. 5. The effect of the variation of photonic parameters is also less visible in the latter. The rapid redistribution and localization towards a few selected states in the latter is a key feature that requires future investigation. It can also be concluded that such a prepared distribution would loose the principal advantages of the entangled photon pair excitations.
@@ III.2.2 Simulations: differential redistribution of two-exciton populations
III.3 Multipoint photon correlation spectroscopy
The quantity of interest in photon correlation spectroscopy is the time-frequency filtered two-photon coincidence counting signal. In this section, we describe how this quantity can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways. The signal can be expressed as:
$S(\{\bar{t}_{j},\bar{\omega}_{j}\})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
@@ III.3 Multipoint photon correlation spectroscopy
where $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}\hat{E}_{sR}^{\dagger}(t^{\prime}+\tau)\hat{E}_{s^{\prime}L}(t^{\prime})$ is a bare photon number operator and $E_{s}(t)=\sqrt{2\pi\hbar\omega_{s}/\Omega}\hat{a}_{s}e^{-i\omega_{s}t}$, and $\Omega$ is the mode quantization volume corresponding to the detector modes. We also defined two detector spectrogram functions: $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$ that are governed by the parameters $\bar{t}_{j}$ and $\bar{\omega}_{j}$, describing the detector centering times for the temporal, and spectral filter, respectively. It can be expressed in terms of the filtering functions as,
@@
Modulation of these functions individually allows for the examination of the photon emission pathways and the systematic recovery of the photon correlation functions. For this work, we choose the Lorentzian filtering functions (see Appendix D), which leads to the spectrogram being expressed as:
@@
We lay out the filtering strategy: the spectral filters allow state-resolution for the dynamics occurring within similar transport windows; the temporal filters help separate the transport processes that involve transitions with similar energy gaps; combined together, they allow for the tracking of individual state transitions. The filtering properties are controlled by the width of the temporal (spectral) filters, $\sigma_{\bar{\omega}_{j}}(\sigma_{\bar{t}_{j}})$, centering times of the temporal (spectral) gate, $\bar{t}_{j}$ ($\bar{\omega}_{j}$). The final signal can be expressed as the sum of four pathways (which is reduced to two following symmetry arguments):
$S(\{\bar{t}_{j},\bar{\omega}_{j}\})$ $=\sum_{p=1}^{4}S^{(p)}(\{\bar{t}_{j},\bar{\omega}_{j}\})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\{t_{j},\omega_{j}\})$
$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
@@
$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
@@
The terms $D_{>}$ and $D_{<}$ indicate the sign of the delay $\tau$ variables, and $\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})$ is the density of states for the detector, taken at an idealized, narrowband limit around each frequency $\omega_{j}$. In our simulation, the influence of this delay was accounted for by evaluating the action of the transport Green’s functions on the initial states within the respective manifolds.
The rationale for deploying such a scheme can be developed in the following manner. During population transport, the ensuing redistribution erases the signatures of the initially created two-exciton wavepacket. The timescales of this phenomena depend on the properties of the transport Green’s functions. The redistribution give rise to new inter-manifold dipole-allowed transitions and render some transitions diminished in spectral intensity.
Figure 7: The time-frequency filtered two-photon coincidence counting signal is displayed, showing variations in the widths of both filtering functions and the centering time of the temporal filtering function. To control the signal resolution, the following systematic variations relative to the reference simulation (upper-left) are presented: the width of the spectral filter (upper-right); the centering time of the temporal filtering function (middle-left); both the spectral filter width and the temporal filter centering time (middle-right); the width of the temporal filter (bottom-left); and the centering time of the temporal filtering function (bottom-right). The targeted excitation of $f_{07}$ by the entangled photon pairs serves as the initial configuration, which is followed by a subsequent series of events: two-exciton manifold transport, inter-manifold transitions, one-exciton manifold transport, and subsequent inter-manifold transitions. For a diagrammatic illustration of the processes involved, see Fig. 2, and for detailed discussion, refer to Section III.3.1
The subsequent photon emission events can be monitored by appropriately selecting the spectral and temporal filtering properties of the first set of detectors. Following the photon emission, the dynamics continue in the one-exciton manifold where transport population scrambling takes place for a time interval, followed by another photon emission event to the ground state. Three representative pathways in Fig.1 shows complexities that may be present.
The joint spectral-temporal filtering strategy can be a powerful tool that enhances the resolution of the bare photon coincidence counting signal. Two sets of spectral-temporal filters, defined by the spectrogram can accomplish two distinct tasks: first, the photon emissions events can be temporally windowed to focus on transport timescales and spectrally windowed to focus on the relevant exciton states; second, the photon emissions events can be spectrally windowed such that pairs of inter-manifold transitions originating from the similar two-exciton transport events can be monitored. Monitoring these events allows us to construct a kinetic description of dissipative exciton dynamics events triggered by the entangled photon excitation. It also help infer how long the remnants of the initially created wavepacket persist. The diagrams presented in Figure 2 clarify the physical processes in a more transparent manner.

### v2

@@ III Results
The sequence of processes: entangled photon excitation, dissipative evolution, and two-photon emission occurs in three stages, as depicted in Fig. 1 (A) (I) to (III). Excitation by entangled photon pairs creates a tailored two-exciton population distribution which undergoes transport for a finite time. It results in a redistribution of populations. The redistributed population serves as the initial configuration for the cascaded two-photon emission processes. The emission process involves two successive one-photon emission events interleaved by one-exciton transport in the one-exciton manifold.
In Fig. 1 (B), we demonstrate three plausible pathways via which the dynamics may proceed. Due to transport, the two-exciton population distribution may lose its narrowband character or may even become narrower. The dynamical information is encoded into the photon emissions that correspond to the inter-manifold transitions between the upper two manifolds ($|f\rangle\rightarrow|e\rangle$). The one-exciton transport that follows may also create additional dispersion of the one-exciton population. The corresponding dynamical information is encoded into the inter-manifold transition between lower two manifolds ($|e\rangle\rightarrow|g\rangle$). These two sets of photon emissions can be classified by controlling the properties of the time and frequency filters, denoted $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$ in the Fig. 1.
In the subsequent sections, we use the following notation for Liouville-space operators: $A_{\alpha}$, where the subscript denotes $\alpha=L,R,+,$-. These operators are defined by their action on a Hilbert-space operator $X$ as $A_{L}X\equiv AX$ and $A_{R}X\equiv XA$. The matrix elements are obtained as $(A_{\alpha}X)_{ij}\equiv\sum_{kl}(A_{\alpha})_{ij,kl}X_{kl}$. We also define $A_{+}\equiv\frac{1}{2}(A_{L}+A_{R})$ and $A_{-}\equiv A_{L}-A_{R}$, where the matrix elements are given by: $(A_{-})_{ij,kl}=A_{ik}\delta_{jl}-A_{\ell j}\delta_{ik}$ and $(A_{+})_{ij,kl}=\frac{1}{2}[A_{ik}\delta_{jl}+A_{lj}\delta_{ik}]$.
The time-frequency filtered two-photon coincidence signal can be expressed as follows:
@@
Here we define the superoperator for filtered detection:
@@
and for the bare two-photon signal
@@
We also use the following superoperators to define the emission processes:
@@
Here the superoperators $T_{L}$ and $T_{R}$ are defined by their action from the left and right sides of the density operator, which describes the initial configuration for the photon emission following the excitation and two-exciton transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$. For future convenience, we also define the observable that corresponds to the preparation of the two-exciton population distribution
$O(t)=\tilde{T}_{R}^{{\dagger}}(\tau_{2},\tau_{3})\tilde{T}_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
@@
Going forward, we assume that the excitation and emission stages are temporally non-overlapping, which results in the factorization of Eq. 7 into separate contributions from each of the three stages. In the following sections, we systematically analyze the role of each stage.
@@ III.1 Entangled photon-pair excitation
The quantity of interest at the preparation stage, the two-exciton population distribution, is defined by the projection operator $|f\rangle\langle f|$ acting on the two-exciton density operator, expressed as $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]$, where the partial trace $\text{Tr}_{B}$ is taken over the phonon modes. We perturbatively expand the density operator to the fourth order in $H_{\text{int}}(t)$ to obtain the expression for the observable:
@@
$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})$
@@
$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (12)
The derivation is outlined in Appendix C. The expression above is a sum of five distinct terms, each representing an exciton pathway that contributes to the signal. The Liouville-space Feynman diagrams in Fig. 3 (I)–(V), offer further insights into the exciton processes captured by this expression. The pathways include contributions from both exciton coherences and populations. The first term represents the two-exciton coherence-mediated pathway (diagram (I)), which is susceptible to inter-manifold dephasing. The second and fourth terms describe pathways susceptible to one-exciton population transport (diagrams (II) and (IV)). The third and fifth terms describe pathways susceptible to intra-manifold dephasing (diagrams (III) and (V)).
The convolutional form provided above is particularly advantageous for exploring the role of entangled photon pairs. Their effect is included via the four-point correlation functions, defined in Eq. II. The following two sections use both degenerate and nondegenerate entangled photon pairs to illustrate these capabilities.
@@
In this section, we monitor the final two-exciton population distribution following excitation by spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. For an SPDC source pumped by a finite-bandwidth laser, the generated photon pairs usually carry a range of frequencies. The photon pairs, emitted within a specific solid angle, must be filtered to select the degenerate components. We carry out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig. 4. The results presented in the left column use an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, which acts as a reference case. The middle and right columns present simulations with a variation in the temporal width of the SPDC pump ($\tau_{0}=50\,\mathrm{fs}$) and the entanglement time parameter ($T_{\mathrm{ent}}=60\,\mathrm{fs}$), respectively, while keeping all other parameters unchanged.
We note that in all three cases, when the target states are located near the middle of the two-exciton manifold, the results exhibit broader population distributions around the target. In other words, we find a lower degree of selectivity. In comparison, when the target states are located in the lower-energy sector of the two-exciton manifold, we observe a higher degree of selectivity. Altering the temporal width of the SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways. In contrast, altering the entanglement time in the right column preserved the selectivity for the majority of states. Across all plots, we notice a trend: some two-exciton states remain preferentially populated across all parameter regimes. This feature has likely emerged from an intricate interplay of exciton transport parameters of mediating states.
@@
We focus on two specific two-exciton states among the full suite of target states previously monitored, namely, $f_{07}$ and $f_{83}$. These two-exciton states reside in the lower ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and the higher ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) energy regions of the two-exciton manifold, respectively. We examine the possibility of exciting these target states via a fixed pair of one-exciton states, selected by setting $\omega_{1}=E_{e_{07}}=15307.1\,\text{cm}^{-1}$ and $\omega_{2}=E_{e_{09}}=15655.9\,\text{cm}^{-1}$. These mediating states are delocalized over multiple chromophoric sites and are prone to ultrafast transport.
These results are displayed in the left columns of Fig. 5 and Fig. 6, respectively. The entanglement time and temporal width of the SPDC pump are set to $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$ for the upper panel, which is similar to the reference case in Fig. 4. In the middle panel, the temporal width is changed to $\tau_{0}=50\,\mathrm{fs}$, whereas in the bottom panel, the entanglement time is changed to $T_{\mathrm{ent}}=40\,\mathrm{fs}$.
The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy in some cases. The increase in the temporal width results in a significant loss of selectivity as the number of mediating pathways with comparable spectral weights grows. The variation of the entanglement time didn’t result in any major changes, except for the tailoring of the relative populations. The latter, however, will have consequences as the population redistributes over time, as will be shown in the next section. We emphasize that both the target states are excited via the same mediating stating one-exciton states. Therefore the differential nature of the final population distribution is dependent on the transition dipoles that connects chosen one-exciton states to the two-exciton target states. The final population distribution demonstrates that the short entanglement time reaches a saturation, less so for the $f_{07}$ than $f_{83}$. The higher lying transitions for the latter imposes stringent requirements on the frequencies of the entangled photon pairs.
@@ III.2 Transport in the two-exciton manifold
The two-exciton population distribution evolves in time, governed by the properties of the transport matrix. The following questions merit further investigation: How does the initial two-exciton population distribution redistribute itself over time? To what extent the two different population distributions differ after undergoing redistribution? Finally, to what extent does the redistribution dynamics depend on the properties of the entangled photon source?
The kinetic equation that governs the dissipative population dynamics of the two-exciton populations is given by
@@
where the transport matrix, $K_{ff,f^{\prime}f^{\prime}}$, follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$. The solution to this equation yields the transport Green’s function $\mathcal{G}_{ff,f^{\prime}f^{\prime}}(\tau)$. The initial two-exciton population distribution, obtained via Eq. III.1, is propagated using the following expression: $\rho_{ff}(t_{2})=\mathcal{G}_{ff,f^{\prime}f^{\prime}}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$. The transport Green’s function is given by the first term in Eq. II and is discussed further in Appendix A.
The following four sets of simulations will help us address the questions posed above.
@@
We monitor the nature of differential population evolution after the excitation targeted at $f_{07}$ (right column in Fig. 5) and $f_{83}$ (right column in Fig. 6), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively. Each of the three panels, upper, middle and bottom, represents simulation using different photonic parameters. The parameters were listed in the preceding section III.1.2 where we discussed the reference simulation. We observe that the evolved distributions at short times retained the memory of prepared distribution in Fig. 5, but gets washed out in Fig. 6. The effect of the variation of photonic parameters is also less visible in the latter. The rapid redistribution and localization towards a few selected states in the latter is a key feature that requires future investigation. It can also be concluded that such a prepared distribution would loose the principal advantages of the entangled photon pair excitations.
@@ III.2.2 Simulations: differential redistribution of two-exciton populations
III.3 Photon correlation spectroscopy
The quantity of interest in photon correlation spectroscopy is the time–frequency filtered two-photon counting signal. In this section, we describe how it can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways.
The rationale for deploying a filtering scheme can be developed as follows. Population transport in the two-exciton manifold acts to redistribute the initially created population. The timescale of this phenomenon is determined by the duration of transport. The population redistribution has two principal consequences: it gives rise to new set of inter-manifold pathways and reduces the spectral intensity of some transitions. Each photon emission pathway can be associated with a specific time and frequency window, which can be selectively monitored by tuning the properties of the first set of time and frequency filters. Following photon emission, the dynamics continue in the one-exciton manifold, where transport may scramble the population distribution. After a finite duration, another inter-manifold transition leads to the ground state. Similar to the above, these transitions can also be monitored via appropriate filtering. Three pathways in Fig. 1 (B) show the dynamical complexities accompany each photon emission. In systems with exciton manifolds, multiple photon emission pathways are possible.
The signal is expressed as:
$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
@@ III.3 Photon correlation spectroscopy
Here, $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}E_{sR}^{\dagger}(t^{\prime}+\tau)E_{s^{\prime}L}(t^{\prime})$ is a photon number operator without the filter. The field operators can be expanded as: $E_{s}(t)=\sqrt{2\pi\omega_{s}/\Omega_{s}}\hat{a}_{s}e^{-i\omega_{s}t}$, where $\Omega$ is the mode quantization volume for the photonic modes of the detector. The latter follows the Bosonic commutation relation $a_{s},a_{s^{\prime}}^{\dagger}]=\delta_{ss^{\prime}}$. We define two spectrogram functions, $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$, which are governed by the time and frequency filtering functions, $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$. The spectrogram function can be represented as
@@
We choose the following time and frequency filters: $F_{\bar{t}}(t^{\prime})=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ and $F_{\bar{\omega}}(\omega^{\prime})=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$. The spectrogram function can be evaluated as
@@
It depends on four parameters: the centering times and frequencies of the time and frequency filters, denoted $\bar{t}_{j}$ and $\bar{\omega}_{j}$, and their respective widths, $\sigma_{\bar{t}_{j}}$ and $\sigma_{\bar{\omega}_{j}}$. One can modulate the spectrogram by varying these parameters. The spectral filters resolve inter-manifold transitions that originate from exciton transport occurring within similar time windows. The temporal filters help separate the transport processes that may involve similar spectral gaps. Combined, they monitor the inter-manifold transitions jointly in time and frequency space.
Using the expression of the spectrogram and following the steps described in Appendix E, we can express the signal as a sum of four pathways (which is reduced to two following symmetry arguments)
$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\sum_{p=1}^{4}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})$
$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
@@
$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
@@
In the above, subscripts in the expression for the spectrograms, $D_{>}$ and $D_{<}$, indicate the sign of the delay $\tau$ between two emissions. The density of photonic states of the detector is assumed to be concentrated at frequencies corresponding to the centering frequency of the frequency filter; these are denoted as $\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})$. In Fig. 3 (I)–(IV), we diagrammatically illustrate the emission pathways in Liouville space. We note that each diagram corresponds to pathways that involve transport in both manifolds, duration of which is controlled by the centering times of the time filters.
Figure 8: The time-frequency filtered two-photon counting signal is displayed. The initial initial population distribution for the photon counting is given by the targeted excitation of $f_{07}$ followed by two-exciton transport. Two sets of time and frequency filters can allow monitoring of the subsequent processes such as inter-manifold transition, one-exciton transport, and inter-manifold transition. The reference simulation in the upper-left uses the following parameters: width of the frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, width of the time filters $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$. In upper-right the the width of the frequency filter is varied to $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$. In middle-left the centering time of the time filter is varied to capture $t_{w,2}=1000\text{fs}$, whereas in middle-right we vary widths of frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ while keeping the and the centering time of the time filter same as middle-left; In bottom-left, the width of the time-filter is varied to $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$. In bottom-right, the centering time of the time filter is varied to capture $t_{w,1}=50\text{fs}$. For detailed discussion, see Section III.3.1